### v1

# Visualizing Thought: Conceptual Diagrams Enable Robust Planning in LMMs

Abstract
Human reasoning relies on constructing and manipulating mental models—simplified internal representations of situations that we use to understand and solve problems. Conceptual diagrams (e.g., a sketch drawn by a human to aid reasoning) externalize these mental models, abstracting irrelevant details to efficiently capture relational and spatial information. In contrast, Large Language Models (LLMs) and Large Multimodal Models (LMMs), predominantly reason through textual representations, limiting their effectiveness in complex multi-step combinatorial and planning tasks. In this paper, we propose a zero-shot fully automatic framework which enables LMMs to reason through multiple chains of (self-generated) intermediate conceptual diagrams, significantly enhancing their combinatorial planning capabilities. Our approach does not require any human initialization beyond natural language description of the task. It integrates both textual and diagrammatic reasoning within an optimized graph-of-thought inference framework, enhanced by beam search and depth-wise backtracking. Evaluated on multiple challenging PDDL planning domains, our method substantially improves GPT-4o’s performance (eg., $35.5\%\rightarrow 90.2\%$ in Blocksworld). On more difficult planning domains with solution depths up to 40, our approach outperforms even the o1-preview reasoning model (eg., over 13% improvement in Parking). These results highlight the value of conceptual diagrams as a complementary reasoning medium in LMMs.
11footnotetext: Xero Inc.22footnotetext: BAIR, UC Berkeley33footnotetext: MIT-IBM Watson AI Lab
Figure 1: Our proposed approach. (1) Model generates an initial diagram schema and conceptual diagram for the domain, visually illustrating spatial information using simple shapes and colors. (2) Inference structured as a graph-of-thought with textual and diagrammatic representations for each state. (3) Model creates diagrams for initial and goal states of the instance conditioned on the conceptual diagram from step (1). (4) Generate each child state and its diagram based on its textual description and parent state. (5) Inference optimizations: Backtracking—if all candidate states at a depth fail verification, backtrack to deepest available ancestor; Beam search—rank candidate states by proximity to goal, selecting top $k=4$ for expansion (Sec. 2). (6) Stop when goal state is reached or maximum allowed states are visited. Example diagrams are from the Tetris domain, which requires rearranging tiles on a grid.
1 Introduction
Natural language is a powerful medium for communication, enabling humans to effectively share knowledge and ideas [44, 7, 49]. However, language alone is not an optimal medium for reasoning, as it is inherently linear, sequential, and verbose, making it inefficient for representing complex logical and relational structures [30, 22, 50]. Prior evidence [28, 16, 22, 4] indicates that humans do not rely solely on language when reasoning. Instead, they construct and utilize internal mental models—simplified analogues of real or hypothetical situations. Humans dynamically manipulate these mental models through mental simulations, which involve representing and predicting interactions between objects, to solve problems. Crucially, mental models are multimodal, integrating both visual and verbal representations to facilitate learning and robust reasoning [37]. Cognitive science research demonstrates that ‘human thought’ is not inherently verbal, sequential, or linear; rather, it is spatial, parallel, and image-like [50]. Finally, visual representations have always played a central role in human reasoning and communication, from prehistoric cave art—one of the earliest known forms of communication, predating written language [8]—to contemporary textbook diagrams, scientific figures in articles, and blackboard sketches.
Sketches, better known as ‘conceptual diagrams’ in the cognitive science literature [30, 50, 22], externalize humans’ internal mental models, drawn to reduce cognitive load and facilitate rapid perceptual inferences and clearer reasoning. Instead of capturing fine-grained details of the external world, similar to the way human eyes perceive the environment, conceptual diagrams abstract away irrelevant details, use simple shapes and colors to represent objects, and emphasize intuitive spatial and relational interactions among entities. [50, 30]. For example, a square in a diagram can represent a complex real-world objects like a car, where the size, color, and position of the shape visually encode its relationships with other entities rather than the entity’s detailed visual appearance. Thus, conceptual diagrams can be an effective reasoning medium complementary to natural language, addressing language’s inherent limitations in accurate representation of spatial and relational structures and aligning closely with humans’ multimodal reasoning [50, 22].
Modern large language models (LLMs) and large multimodel models (LMMs), such as GPT-4 [40] GPT4-o [39] and Gemini [19], have achieved remarkable success on complex mathematical and scientific benchmarks, including GSM8K [9], MATH [24], and GPQA [62]. Despite these advancements, their reasoning performance remains inconsistent, particularly on multi-step compositional reasoning, long-horizon planning, and tasks requiring backtracking and error correction [15, 53, 12]. These limitations partly arise because language models rely solely on language for reasoning, which is inherently inefficient for representing complex relational structures [15, 53]. Additionally, the autoregressive architecture of current LLMs enforces sequential next-token prediction, making tasks that require backtracking particularly challenging [15]. Thus, enabling LLMs to generate and reason with conceptual diagrams within graph-based inference frameworks that enable backtracking is an intuitive and promising approach to alleviate these bottlenecks.
In this work, we propose a framework that enables LMMs to solve combinatorial planning problems by combining textual reasoning with self-generated conceptual diagrams. Our approach operates in a true zero-shot setting, requiring no domain-specific modifications or manual engineering to solve any combinatorial problem that can be expressed in Planning Domain Definition Language (PDDL) format [38], given only the natural language description of the problem (i.e., a clear textual specification of the initial state, goal state, and possible actions). Importantly, our method does not require human specification of the type or style of conceptual diagram to use, and is training-free. Our method works by decomposing inference into a structured graph of intermediate reasoning steps, where at each node the model determines the next best action to take from the current state. Specifically, each node in the inference graph is multimodal, containing both the textual description of the current state and a corresponding conceptual diagram visualizing that state. At each step of inference, the LMM: (i) generates the candidate next state; (ii) verifies uniqueness of the action chosen; (iii) produces a diagram schema — a structured set of textual statements specifying each object’s shape, size, location, and status in the diagram; (iv) Our model autonomously generates executable Matplotlib code for the conceptual diagrams, conditioned on the diagram schema. To ensure inference quality without sacrificing generality, we incorporate guardrails such as diagram schema self-reflection consistency checks and local (parent-child) and global validity checks (entire path from the initial state).
Figure 2: We generate an initial conceptual diagram for each domain by sampling multiple diagram codes. An LMM ranks diagrams based on intuitive and accurate visualization of relational and spatial information. The top-ranked diagram’s code serves as reference for generating initial and goal state diagrams for all instances. Example shown is the parking domain, where curbs hold up to two cars, and cars can be movable or blocked (double-parked).
Figure 3: Child state generation pipeline: the LMM selects an action from the parent node, generates a diagram schema and then an executable diagram code, performs self-reflection and verifies that action chosen does not violate any constraints and action path is feasible (example generations are from Blocksworld domain).
Unlike other approaches that have explored augmenting language models with visual representations for compositional reasoning [59, 56, 26], our method generates conceptual diagrams entirely from scratch, without relying on human-provided initial visual demonstration of the domain or instance, and without requiring tuning of model’s visual encodings. Furthermore, rather than producing a single visualization, our approach autonomously generates multiple sequences of intermediate conceptual diagrams, explicitly illustrating how the LMM’s ‘model’ of the problem evolves with each reasoning step. To address the exponential growth of the combinatorial search space [5, 61], we employ beam search, where validated candidate states at each depth pf the inference graph are ranked by the model based on their proximity to the goal state, and only the top $k$ states are expanded further. Additionally, we incorporate a backtracking mechanism to avoid incomplete searches, allowing the model to revisit previously validated nodes if all candidate states at a given depth fail verification.
Evaluations across multiple challenging PDDL planning domains demonstrate that our method substantially enhances GPT-4o’s combinatorial reasoning capabilities. On the widely studied Blocksworld domain, from Valmeekam et al. [52], our approach significantly improves GPT-4o’s accuracy from 35.5% to 90.2%. Importantly, we contribute a new planning benchmark on five additional complex planning domains — Floor Tiles, Parking, Tetris, Elevator, and Barman —with solution depths up to 40, our method consistently and significantly outperforms GPT-4o, which typically fails entirely (e.g., 0% vs. 36.7% in Floor Tiles, 0% vs. 40% in Tetris). Furthermore, our approach consistently surpasses the best-performing reasoning model, o1-preview [42], across all evaluated complex domains (e.g., 26.7% vs. 40% in Tetris), highlighting the effectiveness of conceptual diagrams as a complementary reasoning medium, enabling LMMs to reliably solve challenging combinatorial planning problems that are otherwise beyond their capabilities, without requiring additional training or domain-specific engineering.
Figure 4: Sequence of intermediate state diagrams in the correct chain, from the initial state to the found goal state, for an example instance of five evaluated domains: Blocksworld, Floor Tile, Elevator, Parking, and Tetris.
2 Method
We propose a zero-shot framework (starting from text-only problem instance description) enabling Large Multi-Modal Models (LMMs), eg., GPT-4o, to solve combinatorial planning problems. Our framework integrates textual reasoning with (self-created by the model) intermediate visual diagrams, built upon a structured graph-of-thought inference approach. Combinatorial planning problems   involve finding a valid sequence of actions from an initial state $s_{0}$ to a goal state $s_{g}$, given a finite set of possible actions. Below, we detail our method following the stages illustrated in Fig. 1: starting from a textual description of the problem instance, (1) generating an initial domain-specific conceptual diagram, (2) visual graph-of-thought inference, (3) initial and goal state diagram generation (4) intermediate state generation pipeline, (5) inference optimizations (beam search and backtracking), and (6) search termination. Detailed prompts for all the steps are provided inside our code that is included an supplementary and would be released upon acceptance.
1. Generating Initial Conceptual Diagram for Domain.
See step 1 in Fig. 1. For each new planning domain, the LMM generates a domain-specific conceptual diagram that visually represents objects and their interactions using simple shapes, colors, and spatial arrangements. This is done entirely in a zero-shot manner, without manual engineering or external visual cues beyond the problem’s textual description. We begin by selecting a random instance from the domain and prompting the LMM to generate multiple candidate diagram schemas. The LMM then ranks these schemas based on how clearly they represent object positions and relationships, selecting the top-ranked one. Using this schema and the problem description, we generate multiple executable Matplotlib diagram codes to visualize the domain (Fig. 2). The LMM then ranks the rendered diagrams based on how effectively they visualize the spatial structure and relationships between objects. The highest-ranked diagram is chosen as the ‘example’ domain-specific conceptual diagram, and its code serves as a reference for generating the diagram code for the initial and goal states of all instances in the domain.
2. Visual Graph-of-Thought Inference.
See step 2 in Fig. 1. Our inference procedure decomposes reasoning into discrete steps, each represented as a node in a structured inference graph, inspired by Graph-of-Thought (GoT)  [5]. With slight abuse of notation we denote by $s_{d}$ an intermediate state at depth $d$ of the graph, represented by a combination of: (i) textual description $T(s_{d})$; (ii) a diagram $D(s_{d})$; and (iii) the action path $A_{0:d}$ from the initial state $s_{0}$. We iteratively expand the inference graph in a breadth-first search (BFS)  [10] manner, expanding nodes on the same depth following the beam search strategy explained in (5) below. Through this inference graph, we simultaneously explore multiple chains of intermediate states toward the goal state.
3. Initial and Goal State Diagram Generation.
See step 3 in Fig. 1. We begin the inference process by generating diagrams for the initial state $s_{0}$ and the goal state $s_{g}$ of each instance, conditioned on the domain conceptual diagram code obtained in step 1. The LMM is instructed to strictly adhere to the representation of objects and their statuses as illustrated in the domain’s conceptual diagram, while accurately initializing the objects according to the specific instance in the diagram code.
4. Intermediate Child State Generation Pipeline.
See step 4 in Fig. 1. From each parent state $s_{d}$, we sample $n=4$ child states. W.l.o.g., next we describe the generation of a single such child state $s_{d+1}$ (Fig. 3). At each node, the LMM first selects the next candidate action $a_{d+1}$, conditioned on the parent state $s_{d}$ (which is represented by its diagram, textual description, and action path from initial state). We then generate the textual description $T(s_{d+1})$ of the resulting child state. The candidate action-state pair $(a_{d+1},T(s_{d+1}))$ is compared to previously generated child states to verify uniqueness. If different, the LMM generates a $s_{d+1}$ specific instance of the problem’s conceptual mapping $E(s_{d+1})$, which is a set of textual statements specifying shapes, sizes, positions, statuses (e.g., colors), and textual identifiers for each object in the state (see Fig. 3 for an example). The $E(s_{d+1})$ undergoes a self-reflection verification step to ensure consistency with $T(s_{d+1})$ and accuracy with respect to $a_{d+1}$, if failed we regenerate it. Subsequently, Matplotlib code $C(s_{d+1})$ is generated conditioned on $E(s_{d+1})$, $T(s_{d+1})$, and two example diagram codes: the initial state diagram code $C(s_{0})$ and the parent state diagram code $C(s_{d})$. Code is re-generated if it fails to run. After successfully generating the child state diagram, we perform two additional action verification checks: (1) a local check that verifies the validity of the action $a_{d+1}$ based solely on the diagrams and textual descriptions of the parent and child states (ensuring the action complies with domain constraints); and (2) a global check that verifies whether the entire action path $A_{0:d+1}$ remains feasible for reaching the goal state $s_{g}$. If either of these checks fails, the child state is marked as invalid.
5. Inference Optimizations: Beam Search and Backtracking.
See steps 5 in Fig. 1. Applying Graph-of-Thought [5] to combinatorial planning problems via naively expanding all nodes (eg., standard BFS) would result in exponential growth of the search tree [46]. To mitigate this, we employ two inference optimizations: beam search and depth-wise backtracking. Using a method inspired by beam search  [35], we first expand all states at each depth $d$ using BFS, generating up to $N$ child states per state. The LMM then ranks all candidate child states at depth $d+1$ based on their proximity to the goal state $s_{g}$, selecting only the top $k=4$ states for further expansion. This depth-wise ranking solves the exponential growth problem. Moreover, this ranking system resembles human problem-solving strategies, where shallow state-specific heuristics are employed to estimate how close intermediate states are to the goal  [18, 6, 45]. Additionally, we implement a depth-wise backtracking mechanism. If all candidate child states at a given depth fail verification, we backtrack to the deepest available ancestor nodes at depth $d_{\text{max}}$ and attempt new expansions. We allow up to $B=2$ backtracking attempts to any given depth. If all $B$ attempts at depth $d_{\text{max}}$ fail, we mark nodes at that depth as invalid and backtrack further to the next deepest available nodes at depth $d_{\text{max2}}$, where $d_{\text{max2}}<d_{\text{max}}$.
6. Search Termination.
See step 6 in Fig. 1. The inference process continues iteratively, expanding nodes depth-by-depth, until either the goal state $s_{g}$ is reached or a predefined computational budget is exhausted. Specifically, we set two types of computational limits: (1) a maximum number of generated states (120 states for simpler Blocksworld instances, 450 states for more complex domains), and (2) a maximum depth limit (28 for simpler Blocksworld instances, 100 for other domains). These limits ensure computational efficiency in terms of inference time and API usage costs. If the goal state is not reached within these constraints, the search is marked as incomplete. The goal state verification is performed by the LMM, which compares the diagram and textual description of the current state against the description of the goal state and its diagram (generated in (3) above).
In summary, our method leverages structured visual reasoning, self-generated conceptual diagrams, and optimized graph-based inference strategies to efficiently solve combinatorial planning problems with an LMM. Each step of our pipeline is visually illustrated in Figures 12, and 3.
Models Claude 3.5 (Sonnet) Claude 3 (Opus) GPT-4o GPT-4 Turbo LLaMA 3.1 405B Gemini 1.5 Pro o1-preview o1-mini GPT4-o + Visual Thinking One Shot 57.6% 48.1% 28.3%* 23% 47.3% 16.8% - - - Zero Shot 54.8% 59.3% 35.5% 40.1% 62.6% 23.8% 97.8% 56.6% 90.2%
Table 1: Performance of different models in the Blocksworld domain. Baseline numbers taken from [54]
3 Related Work
Multimodal Chain-of-Thought for Reasoning. Several recent works have explored integrating visual representations into reasoning processes for LLMs and LMMs. [26] equips multimodal models with drawing tools to generate intermediate sketches composed of lines and shapes, but primarily addresses single-step problems with only a single intermediate visualization per problem instance. Similarly, [56] generates visual aids for spatial reasoning, providing a single refined visualization per problem, which typically is an approximate of a high-fidelity illustration rather than a conceptual diagrams representing object interactions. Concurrent work, [59], generates visual and textual intermediate steps for planning tasks by conditioning multi-modal models on externally provided initial visual representations that supplement textual problem descriptions. Their approach iteratively generates intermediate textual and visual states, but the reliance on externally supplied initial visualizations could limit generality and applicability to unseen domains.
Our approach differs in several key aspects. First, our method autonomously generates the first conceptual diagram directly from textual problem descriptions, without relying on external visual demonstrations or cues. Second, rather than producing a single visualization, our approach creates multiple chains of intermediate visual states, enabling parallel multi-hypothesis compositional reasoning through evolving diagrams. Third, our diagrams explicitly represent relationships between objects in a conceptual manner, rather than realistic or approximate visual depictions of the objects. Finally, our method operates in a zero-shot setting, capable of solving any problem expressible in PDDL format (very general, compatible with most planning problems) without prior visual examples or domain-specific visual guidance. These distinctions collectively enable a more generalizable and flexible form of diagrammatic reasoning, leading to significant performance gains.
Search and Verification Strategies for Enhanced LLM Reasoning. Recent methods improve reasoning in LLMs by structuring inference into explicit intermediate steps [58, 29] and employing search-based strategies over multiple reasoning paths [57, 61, 21, 11, 5]. For instance, Yao et al. [61] propose Tree-of-Thoughts (ToT), which extends CoT to explore a tree of reasoning paths, and Besta et al. [5] introduce Graph-of-Thoughts (GoT), which structures the reasoning process as a graph, enabling exploration and aggregation of intermediate reasoning steps. Other works utilize verification and refinement techniques, such as iterative self-reflection and feedback, to enhance reasoning accuracy [36, 43, 48]. Our approach builds upon these methods by integrating conceptual diagrams into intermediate reasoning steps. Moreover, as detailed in Section 2, we extend graph-based inference with beam search and depth-wise backtracking, significantly reducing the search space and improving performance on long-horizon planning tasks.
World Modeling and Planning with Language Models.. Recent research explores planning and reasoning using LLMs by implicitly or explicitly constructing world models from textual descriptions [27, 60, 3, 13, 14, 33, 55]. For instance, Huang et al. [27] show that LLMs implicitly form textual world models for simple planning tasks, while other works explicitly represent states and transitions for structured reasoning [60, 13, 33]. Recent approaches also introduce explicit state-tracking, modeling entity relationships and transitions to enhance reasoning [33, 55, 14]. Our work extends these approaches by enabling LMMs to autonomously construct and reason with diagrammatic world models. Specifically, our method generates conceptual diagrams directly from textual problem descriptions, defining a visual schema of objects and their relationships through abstract shapes, colors, and spatial relationships. Subsequent diagram updates visually simulate action sequences, explicitly tracking object states and interactions. Compared to purely textual world modeling, our diagrammatic approach provides clearer and more accurate representations of spatial and relational information, significantly improving performance on complex, multi-step planning tasks.
Domain GPT4o o1-mini o1-preview GPT4-o + Visual Thinking
Blocksworld (hard) 0% 10% 23.65% 78%
Floor Tiles 0% 6.7% 23.3% 36.7%
Parking 3.3% 6.7% 36.7% 50%
Tetris 0% 3.33% 26.67% 40%
Elevator 0% 3.33% 33.3% 46.67%
Barman 0% 0% 13.3% 33.3%
Table 2: Performance of select models on various domains.
4 Evaluation
To evaluate the proposed approach we have conducted experiments 7 different planning problem domains, including the popular Blocksworld [20] and Blocksworld Hard [54]; as well as 5 new problem domains contributed in this work, prepared by us using a popular repository of PDDL planning problem domains [1].
4.1 Implementation Details
We evaluated our method using GPT-4o as the LMM, and o1-preview [42] and o1-mini [41] as baselines, all accessed via the OpenAI API. Baseline models solved instances in PDDL format using prompts from [54]. Plan correctness was determined with VAL [25]. Experiments ran on a machine with 96 Intel Xeon 5220R CPUs (4GHz, 71.5 MiB L3 cache). On this machine, processing each depth level took approximately 1 minute. Consequently, simpler Blocksworld instances typically took around 5–8 minutes per instance, which is comparable to the runtime of the o1-preview model. Our code is provided in the supplementary and would be released upon acceptance.
Analysis Corrects Incorrects Incompletes Avg Depth of Corrects Max Depth of Corrects Min Depth of Corrects Avg Number of States Generated
Blocksworld (hard) 78% 10% 12% 20.15 36 18 177.7
Floor Tiles 36.7% 23.3% 40% 15.88 25 10 244
Parking 50% 40% 10% 9.6 29 2 149.1
Tetris 40% 60% 0% 6.45 11 4 46.24
Elevator 46.67% 43.33% 10% 20.89 26 16 160.30
Barman 33.33% 40% 26.67% 24.0 27 22 210.54
Table 3: Analysis of domains. ‘Corrects’: % instances with correct plans; ‘Incorrects’: % with incorrect plans; ‘Incompletes’: % terminated due to state budget; ‘Avg/Max/Min Depth of Corrects’: avg/max/min number of actions in correct plans; ‘Avg Number of States Generated’: avg states generated across all instance of the domain. ‘Corrects’ is our main metric.
Ablation Corrects Incorrects Incompletes Avg Depth of Corrects Max Depth of Corrects Avg Number of States Generated
GPT4-o + Visual Thinking 90.2% 6.4% 3.4% 10.07 28 38.89
No Diagram 58% 36% 6% 8.28 18 25.86
No Diagram Schema 72% 20% 8% 7.41 18 46.2
No Code Execution 24% 66% 10% 7.82 22 32.7
1-Branching Factor 52% 4% 44% 6.38 16 23.9
2-Branching Factor 70% 6% 24% 8.29 24 26.27
No Backtracking 62% 4% 34% 6.06 14 22.12
No Beam Search 72% 4% 24% 6.61 12 58.31
Table 4: Ablation Study Results. ‘Corrects’ is our main metric. See Tab. 3 for columns notation.
4.2 Datasets
Evaluated Domains.
We evaluate our method on combinatorial planning problems [17] from the International Planning Competition (IPC) [2], expressed in PDDL format [38]. These problems are mostly NP-hard [20]. We evaluated out method on six domains, namely Blocksworld [20], and 5 other IPC domains: Floor Tile [23], Parking, Tetris [51], Elevator, and Barmen. For each domain, we generated instances using publicly available instance generators [1]. The Blocksworld domain involves stacking and unstacking blocks. We used the 500 simple instances in PlanBench [52] with 3-5 blocks and 30 additional hard instances with 10-20 blocks (same instance complexity as in [54]). The Floor Tile domain features robots painting tiles on a grid. We generated 30 PDDL instances for this domain with varying grid sizes (2-3 rows, 3-5 columns) and 1-2 robots. The Parking domain involves rearranging cars in curbs, with 30 instances using 4-5 curbs and 4-6 cars. The Tetris domain requires rearranging Tetris tiles on a grid, with 30 instances using $(4\times 4)$ or $(6\times 6)$ grids. Lastly, the Elevator domain simulates passenger transport in buildings, with 30 instances using 4-5 floors and 10-12 passengers. Refer to Fig. 4 a sequence of intermediate state diagrams in the correct chain for an example instance of a subset of evaluated domains.
Translating PDDL to Natural Language and Back.
Our method operates entirely over natural language descriptions. Thus, we first translate PDDL domains and instances into natural language. For domain translation, we use a manually engineered five-shot prompt containing translations of five different PDDL domains into natural language. For instance translation, we use a one-shot prompt containing a manually translated instance from the same domain. After solving the problem using our pipeline, we translate the resulting natural language action sequence back into PDDL using a one-shot prompt, which includes a random partial plan from the same domain along with its syntax-accurate translation to PDDL.
4.3 Results
Our main results are summarized in Tables 1 and 2. Table 1 compares the performance of our method (using GPT-4o as the LMM) on the popular, yet simpler Blocksworld planning domain  against leading closed- and open-source models at the time of writing, including the specialized reasoning models o1-preview and o1-mini. Most baseline results are taken from [54]. Our approach significantly outperforms the base GPT-4o model by approximately 55%, achieving 90.2% accuracy, and outperfoms all baselines, except for a slight disadvantage compared to the leading reasoning model, o1-preview (97.8%). We conjecture two reasons for o1-preview’s advantage: (i) the smaller problem size and shallow solution depth of simpler Blocksworld instances, and (ii) This planning domain is widely used, with many solved instances publicly available online, likely providing extensive exposure to o1 during its training.
This conjecture is supported by our results in Table 2, which evaluates performance on significantly larger and deeper Blocksworld (hard) instances, as well as five additional challenging domains derived from the IPC competition [2], where new harder instances are generated and solutions are not publicly available. In these more complex and unfamiliar domains, our method substantially outperforms all baselines, including o1-preview. For instance, in Blocksworld (hard), our method achieves 78% accuracy compared to only 23.65% by o1-preview. Similarly, our method consistently surpasses o1-preview across Floor Tiles, Parking, Tetris, Elevator, and Barman domains, demonstrating the robustness and generalizability of our structured visual reasoning approach.
Table 3 provides further insights into our method’s performance on these challenging domains. Despite significantly deeper solution paths (with average depths up to 24 actions per instance and up to 244 states generated per instance), our method successfully solves instances requiring as many as 36 sequential actions (Blocksworld hard). In contrast, GPT-4o alone achieves very low performance (less than 0-4%), highlighting its susceptibility to compounding errors in long-horizon compositional reasoning [54, 15]. Our structured visual reasoning approach, incorporating backtracking and verification steps, effectively mitigates these compounding errors, enabling more robust performance even in deep and complex planning problems.
The primary limitation of our method observed in these experiments is the number of incomplete searches (e.g., 40% incomplete in Floor Tiles, 26.67% in Barman) which arise due to the computational budget constraint set. Additionally, we note the method exhibits the highest incorrect rate (60%) on Tetris domain, primarily due to its inherently high branching factor (up to 24 possible actions per state) and complex action parameterization—each action can require up to 7 parameters detailing positions of all sub-tiles, compared to simpler domains like Blocksworld, where actions typically require only 1-2 parameters. Conversely, our largest performance gain over o1-preview occurs in the Barman domain (33.3% vs. 13.3%), likely due to our diagrammatic representation efficiently capturing complex object relationships and high number of interactions, which are challenging to represent for purely textual reasoning models.
4.4 Ablation Studies
To systematically evaluate the contributions of different components of our proposed framework, we conducted ablation studies on 50 instances from the Blocksworld domain. We examined the impact of state diagrams, diagram schema, diagram code execution, different branching factors, and inference optimizations (beam search and backtracking). Table 4 summarizes the results of these experiments.
Impact of Various Components of State Diagram Generation.
We first assessed the effectiveness of state diagrams by removing them entirely from the inference pipeline (”No Diagram”). This resulted in a substantial performance drop from 90.2% to 58%, highlighting the critical role diagrams play in succinctly representing spatial and relational information. Additionally, the average solution depth of correctly solved instances decreased from 10.07 to 8.28, indicating that without diagrams, the model struggled to solve more complex instances requiring deeper reasoning.
In another experiment, we removed the diagram schema (”No Diagram Schema”) from the child state generation pipeline, directly inferring diagram code from the textual description of the state. This caused performance to drop from 90.2% to 72%. Although less severe than removing diagrams entirely, this drop demonstrates that diagram schemas effectively highlight spatial information from the textual description, aiding accurate diagram generation.
Finally, we tested removing the rendered diagrams, providing only the executable Matplotlib code (”No Code Execution”) when generating the next action. This resulted in the most significant performance drop (90.2% to 24%), even below the GPT-4 baseline (36%), clearly demonstrating that using the diagram code directly distracts the model and impairs reasoning. This also aligns with prior findings that extraneous details negatively impact model performance [47, 34]. These results reinforce the importance of diagrams as compact, intuitive representations that facilitate rapid perceptual inference and clear spatial reasoning [30, 50, 22]. Even though diagrams encode the same information as textual descriptions or code, it is the representation of information in a two-dimensional plane which significantly aids spatial understanding and inference.
Impact of Branching Factor.
We next investigated the effect of branching factor of the inference graph (the number of candidate child states generated per state) on performance. Reducing the branching factor from 4 (our default) to 2 (”2-Branching Factor”) decreased performance from 90.2% to 70%, primarily due to a significant increase in incomplete searches (24% vs. 3.4%). This suggests that the first two candidate states, that in our approach are generated at lower sampling temperatures, include less correct next states, and exploring the third or fourth candidates, generated at higher temperatures, is frequently necessary to find the correct solution path. Thus, reducing branching factor limits randomness and diversity in candidate generation, leading to more incomplete searches. Additionally, the average depth of correctly solved instances decreased from 10.07 to 8.29, indicating difficulty solving problems with deeper solution depths with fewer candidate states. Further reducing the branching factor to 1 (”1-Branching Factor”), effectively converting our graph-based inference into a chain-of-thought approach (with diagrams), caused performance to drop even further to 52%. Despite this significant drop, performance remained substantially higher than the GPT-4 baseline (36%), underscoring the value of diagrams in improving reasoning performance of the LMM even without extensive search.
Impact of Inference Optimizations (Backtracking and Beam Search).
Finally, we evaluated the importance of our inference optimizations (backtracking and beam search) on top of the graph-of-thought (with diagrams) approach. Removing backtracking (”No Backtracking”) significantly reduced performance from 90.2% to 62%, primarily due to a large increase in incomplete searches (34% vs. 3.4%). This increase arises because the LMM verification steps occasionally produce false negatives, incorrectly invalidating correct states. Without backtracking, the model cannot recover from these errors, leading to incomplete searches as no valid nodes remain at the frontier of the inference graph to expand from. Similarly, removing beam search (”No Beam Search”) reduced performance from 90.2% to 72%, again primarily due to increased incompletes (24% vs. 3.4%). However, unlike backtracking, incompletes here result from exponential growth in the search space, causing the model to exhaust its computational budget (maximum allowed number of states to generate) before reaching the goal state. Indeed, the average number of states generated increased significantly (58.31 vs. 38.89), highlighting beam search’s critical role in efficiently managing combinatorial explosion. Both optimizations proved essential for solving deeper combinatorial problems, as evidenced by the reduced average depth of solved instances (6.61 without beam search, 6.06 without backtracking, vs. 10.07 of our full approach). These results demonstrate that backtracking and beam search are complementary and crucial optimizations for effectively solving complex combinatorial planning problems using graph-based inference.
5 Conclusion
Contributions.
In this paper, we introduced a zero-shot framework enabling LMMs to robustly solve combinatorial planning problems by integrating textual reasoning with self-generated conceptual diagrams. Our key contributions include: (1) autonomously generating conceptual diagrams directly from textual descriptions; (2) structuring inference as a visual graph-of-thought with multiple chains of intermediate textual and diagrammatic representations; (3) employing beam search and depth-wise backtracking for efficient combinatorial search; and (4) significantly outperforming GPT-4o and the specialized reasoning model o1-preview across multiple challenging planning domains.
Future Work.
Besides enhancing reasoning, our generated conceptual diagrams can directly benefit downstream tasks such as grounded video and image generation, where they can be used as explicit scene layouts which have shown to significantly improve quality [32, 31]. Our structured inference naturally produces accurate diagrams grounded in provided natural language constraints, offering a promising pathway toward fully automatic visual generation for complex, long-horizon tasks.
Supplementary Material