### v2

# Visualizing Thought: Conceptual Diagrams Enable Robust Combinatorial Planning in LMMs

Abstract
Human reasoning relies on constructing and manipulating mental models—simplified internal representations of situations that we use to understand and solve problems. Conceptual diagrams (e.g., a sketch drawn by a human to aid reasoning) externalize these mental models, abstracting irrelevant details to efficiently capture how entities interact with each other. In contrast, Large Language Models (LLMs) and Large MultiModal Models (LMMs) predominantly reason through text, limiting their effectiveness in complex multi-step tasks. In this paper, we propose Visual Thinking, a zero-shot framework that enables LMMs to reason through multiple chains of (self-generated) conceptual diagrams, significantly enhancing their combinatorial planning capabilities. Our approach does not require any human initialization beyond the natural language description of the task. It integrates both textual and diagrammatic reasoning within an optimized Graph-of-Thought inference framework, enhanced by beam search and depth-wise backtracking. Evaluated on multiple challenging PDDL planning domains, our method substantially improves LMMs’ performance (e.g., GPT-4o: $35.5\%\rightarrow 90.2\%$ in Blocksworld) and consistently outperforms other text-only search-based inference methods. On more difficult planning domains with solution depths up to 40, our approach outperforms even the o1-preview reasoning model (e.g., 16 percentage points improvement in Floor Tiles). These results highlight the value of conceptual diagrams as a reasoning medium in LMMs.
00footnotetext: 1Xero Inc.  2BAIR, UC Berkeley  3MIT-IBM Watson AI Lab
1 Introduction
Natural language is a powerful medium for communication, enabling humans to effectively share knowledge and ideas Pinker and Bloom (1990); Clark (1996); Tomasello (2010). However, language alone is not an optimal medium for reasoning, as it is inherently linear, sequential, and verbose, making it inefficient for representing complex logical and relational structures Larkin and Simon (1987); Hegarty (2004); Tversky (2011). Prior evidence Johnson-Laird (1983); Gentner and Stevens (2001); Hegarty (2004); Battaglia et al. (2013) indicates that humans do not rely solely on language when reasoning. Instead, they construct and utilize internal mental models—simplified analogues of real or hypothetical situations. Humans dynamically manipulate these mental models through mental simulations, which represent and predict interactions between objects, to solve problems. Crucially, mental models are multimodal, integrating both visual and verbal representations to facilitate learning and robust reasoning Mayer (2002). Cognitive science research demonstrates that ‘human thought’ is not inherently verbal, sequential, or linear; rather, it is spatial, parallel, and image-like Tversky (2011). Finally, visual representations have always played a central role in human reasoning and communication, from prehistoric cave art—one of the earliest known forms of communication, predating written language Clottes (2008)—to contemporary textbook diagrams, scientific figures in articles, and blackboard sketches.
Conceptual diagrams are simplified visual representations that use basic shapes to capture how entities interact with each other while abstracting away irrelevant details Tversky (2011); Larkin and Simon (1987). These diagrams externalize the internal mental models of humans, reducing cognitive load and allowing rapid perceptual inference and clearer reasoning Hegarty (2004); Gentner and Stevens (2001). Unlike images, which capture the fine-grained details of how objects appear in the external world as perceived by the human eye, conceptual diagrams emphasize the structural and relational information needed for reasoning: they use simple shapes, colors, and relative positions and sizes to represent interactions among entities Tversky (2011); Larkin and Simon (1987). For example, a square in a diagram can represent a complex real-world object such as a car, where its size, color, and position visually encode its relationships with other entities abstracting away the car’s detailed appearance. Thus, conceptual diagrams are an effective reasoning medium complementary to natural language, addressing language’s inherent limitations in accurate representation of relational structures and aligning closely with humans’ multimodal reasoning Tversky (2011); Hegarty (2004); Dziri et al. (2023).
Modern large language models (LLMs) and large multimodal models (LMMs), such as GPT-4 OpenAI (2023) GPT-4o OpenAI (2024a) and Gemini Google AI (2025), have achieved remarkable success on complex mathematical and scientific benchmarks, including GSM8K Cobbe et al. (2021), MATH Hendrycks et al. (2021), and GPQA Zhong et al. (2023). Despite these advancements, their reasoning performance remains inconsistent, particularly on multi-step compositional reasoning, long-horizon planning, and tasks requiring backtracking and error correction Dziri et al. (2023); Valmeekam et al. (2023b); Creswell et al. (2023). These limitations partly arise because language models rely solely on language for reasoning, which is inherently inefficient for representing complex relational structures Dziri et al. (2023); Valmeekam et al. (2023b). Additionally, the autoregressive architecture of current LLMs enforces sequential next-token prediction, making tasks that require backtracking particularly challenging Dziri et al. (2023). Thus, enabling LLMs to generate and reason with conceptual diagrams within graph-based inference frameworks that enable backtracking is an intuitive and promising approach to alleviate these bottlenecks.
Figure 1: Our proposed approach. Example diagrams are from the Tetris domain, where tiles are moved on a grid to reach a goal state. (1) The model generates multiple diagram schemas and codes for a random instance; their rendered diagrams are ranked; the code of the top choice is cached (Fig. 2). (2) Conditioned on this code, the model produces diagrams for the initial and goal states. (3) Candidate child state, along with its diagrams, is generated based on the parent state (Fig. 3). (4) Beam search ranks all candidates at a depth by proximity to the goal, expanding the top $k=4$ (Sec. 3). (5) Depth-wise backtracking is applied when all candidate states at a depth fail validation, returning to the deepest available ancestor. (6) The process stops when the goal is reached or a maximum number of steps is visited. (7) The output is the action sequence (plan) plus textual and diagrammatic representations of intermediate states.1
11footnotetext: For Figs. 1 and 4, we adjusted the diagrams to increase font sizes for better readability. Figs. 2-3 are unmodified generations. Fig. 1 action descriptions are simplified; full action strings include previous/current cells occupied.
In this work, we propose a framework that enables LMMs to solve combinatorial problems by combining textual reasoning with self-generated conceptual diagrams. Our approach operates in a zero-shot setting, requiring no domain-specific modifications or manual engineering to solve any combinatorial problem that can be expressed in Planning Domain Definition Language (PDDL) format McDermott et al. (1998), given only the natural language description of the problem (i.e., a textual specification of the problem’s initial state, goal state, and possible actions). Importantly, our method does not require any visual templates or geometric specifications; instead, it generates conceptual diagrams entirely from the textual description of the problem.
Our method decomposes inference into a graph of intermediate reasoning steps, where at each node the model determines the next best state from the current state. Each node is multimodal, containing both the textual description of the state and a corresponding conceptual diagram. At each step, the LMM: (i) generates the candidate next state using the textual and diagrammatic representations of previous states; (ii) produces a diagram schema—a structured set of textual statements specifying each object’s shape, relative size and location, and status; and (iii) autonomously generates executable Matplotlib code that renders the conceptual diagram of the state, conditioned on the schema. To ensure inference quality, we incorporate guardrails such as diagram schema self-reflection checks and local (parent-child) and global (entire path from the initial state) validity checks.
Unlike prior approaches that augment language models with visual representations for compositional reasoning Wu et al. (2024); Wang et al. (2024); Hu et al. (2024)—where models are given an initial visual demonstration or template of the state and then asked to update it—our method relies solely on the textual description of each state. The model independently generates conceptual diagrams for every state entirely from scratch, without any human-provided visual examples, templates, or cues, and without any tuning of the model’s visual encodings. This end-to-end diagram generation process mirrors how humans use imagination to construct mental models from language. Furthermore, rather than producing a single visualization, our approach autonomously generates multiple sequences of intermediate conceptual diagrams, explicitly illustrating how the LMM’s ‘model’ of the problem evolves with each reasoning step. To address the exponential growth of the combinatorial search space Besta et al. (2023); Yao et al. (2023), we employ beam search, where validated candidate states at each depth of the inference graph are ranked by the model based on their proximity to the goal state, and only the top $k$ states are expanded further. Additionally, we incorporate a backtracking mechanism to avoid incomplete searches, allowing the model to revisit previously validated nodes if all candidate states at a given depth fail verification.
Evaluations across multiple challenging PDDL planning domains demonstrate that our method substantially enhances LLMs’ combinatorial reasoning capabilities. On the widely studied Blocksworld domain Valmeekam et al. (2023a), our approach yields performance gains of roughly 43, 64, and 55 percentage points using Claude 3.5 Sonnet Anthropic (2024), LLaMA 4 Maverick Meta (2025), and GPT-4o, respectively (e.g., GPT-4o’s accuracy rises from 35.5% to 90.2%). Importantly, we contribute a new planning benchmark on five additional domains—Floor Tiles, Parking, Tetris, Elevator, and Barman—with solution depths designed up to 40, where our method consistently succeeds even as the base model (GPT-4o) fails entirely (e.g., 0% vs. 36% in Floor Tiles, 0% vs. 38% in Tetris). Furthermore, our approach with GPT-4o as the base model, surpasses the best-performing reasoning model, o1-preview OpenAI (2024c), across all evaluated complex domains (e.g., 26% vs. 38% in Tetris). Finally, compared to strong search-based planning methods such as Graph-of-Thought and RAP (a Monte Carlo Tree Search framework that uses an LLM both as world model and planner), our method improves accuracy by at least 22 percentage points while also reducing inference cost by over 30% and latency by more than 25%. These results highlight conceptual diagrams as an efficient reasoning medium, enabling LMMs to reliably solve challenging long-horizon planning problems that remain beyond the reach of purely textual (single inference or search-based) approaches.
Single inference Search-based inference (base LLM = GPT-4o)
Domain
GPT-4o
o1-
mini
o1-
preview
GoT
Optimized
GoT
RAP(20)
Visual Thinking
Blocksworld (simple) 35.5% 56.6% 97.8% 50% 58% 58% 90.2%
Blocksworld (hard) 0% 0.9% 23.65% 8% 48% 4% 78%
Floor Tiles 0% 6% 20% 0% 4% 36%
Parking 2% 8% 40% 14% 28% 52%
Tetris 0% 2% 26% 0% 12% 38%
Elevator 0% 2% 36% 2% 10% 48%
Barman 0% 0% 10% 0% 4% 30%
Table 1: Accuracy results across all evaluated domains. Each baseline was evaluated on 50 problem instances per domain, except for GPT-4o + Visual Thinking on Blocksworld (simple), where we evaluated 500 instances (the full PlanBench size Valmeekam et al. (2023a)) and achieved 90.2% (451/500). Baseline results marked with are taken from Valmeekam et al. (2024). RAP reported 51% accuracy on Blocksworld (simple) using LLaMA 2 Hao et al. (2023).
2 Related Work
Multimodal Chain-of-Thought for Reasoning. Several recent works have explored integrating visual representations into reasoning processes for LLMs and LMMs. Hu et al. (2024) equips multimodal models with drawing tools to generate intermediate sketches composed of lines and shapes, but primarily focuses on single-step or shallow problems, typically producing one intermediate visualization per instance. Similarly, Wang et al. (2024) generates visual aids for spatial reasoning, providing a single refined visualization per problem. In both works, generated visualizations are typically an approximate or an augmentation of a high-fidelity illustration rather than conceptual diagrams representing object interactions. Concurrent work, Wu et al. (2024), generates visual and textual intermediate steps for planning tasks by conditioning multimodal models on externally provided initial visual representations that supplement textual problem descriptions. Their approach iteratively generates intermediate textual and visual states, but the reliance on externally supplied initial visualizations could limit generality and applicability to unseen domains.
Our approach differs in several key aspects. First, our method autonomously generates conceptual diagrams directly from textual problem descriptions, without relying on external visual demonstrations or cues, mirroring the human ability to construct mental models from language. Second, rather than producing a single visualization, our approach creates multiple chains of intermediate visual states, enabling parallel multi-hypothesis compositional reasoning through evolving diagrams. Third, our diagrams explicitly represent relationships between objects in a conceptual manner, rather than realistic or approximate visual depictions of the objects. Finally, our method operates in a zero-shot setting, and is applicable to any problem expressible in PDDL format (a general language for planning problems) without any domain-specific information required. These distinctions collectively enable a more generalizable and flexible form of diagrammatic reasoning, leading to significant performance gains.
Search and Verification Strategies for Enhanced LLM Reasoning. Recent methods improve reasoning in LLMs by structuring inference into explicit intermediate steps Wei et al. (2022); Kojima et al. (2022) and employing search-based strategies over multiple reasoning paths Wang et al. (2022); Yao et al. (2023); Hao et al. (2023); Creswell et al. (2022); Besta et al. (2023). For instance, Yao et al. (2023) propose Tree-of-Thought (ToT), which extends CoT to explore a tree of reasoning paths, and Besta et al. (2023) introduce Graph-of-Thought (GoT), which structures the reasoning process as a graph, enabling exploration and aggregation of intermediate reasoning steps. Other works utilize verification and refinement techniques, such as iterative self-reflection and feedback, to enhance reasoning accuracy Madaan et al. (2023); Paul et al. (2023); Shinn et al. (2023). Our approach builds upon these methods by integrating conceptual diagrams into intermediate reasoning steps. Moreover, as detailed in Section 3, we extend graph-based inference with beam search and depth-wise backtracking, significantly reducing the search space and improving performance on long-horizon planning tasks.
World Modeling and Planning with Language Models. Recent research explores planning and reasoning using LLMs by implicitly or explicitly constructing world models from textual descriptions Huang et al. (2022); Yao et al. (2022); Ahn et al. (2022); Dasgupta et al. (2022); Driess et al. (2023); Liu et al. (2023a); Wang et al. (2023). For instance, Huang et al. (2022) show that LLMs implicitly form textual world models for simple planning tasks, while other works explicitly represent states and transitions for structured reasoning Yao et al. (2022); Dasgupta et al. (2022); Liu et al. (2023a). Recent approaches also introduce explicit state-tracking, modeling entity relationships and transitions to enhance reasoning Liu et al. (2023a); Wang et al. (2023); Driess et al. (2023). Our work extends these approaches by enabling LMMs to autonomously construct and reason with diagrammatic world models. Specifically, our method generates conceptual diagrams directly from textual problem descriptions, defining a visual schema of objects and their relationships through abstract shapes, colors, and spatial relationships. Subsequent diagram updates visually simulate action sequences, explicitly tracking object states and interactions. Compared to purely textual world modeling, our diagrammatic approach provides clearer and more accurate representations of spatial and relational information, significantly improving performance on complex, multi-step planning tasks.
Figure 2: We generate an initial conceptual diagram for each domain by sampling multiple diagram codes. An LMM ranks diagrams based on intuitive and accurate visualization of relational information. The top-ranked diagram’s code serves as reference for generating initial and goal state diagrams for all instances. Example shown is the parking domain, where curbs hold up to two cars, and cars can be movable or blocked (double-parked).
3 Method
We propose Visual Thinking, a training-free and model-agnostic framework, which integrates textual reasoning with model-generated intermediate conceptual diagrams, to enable Large MultiModal Models (LMMs) to solve combinatorial problems, given text-only problem specifications (initial state, goal, admissible actions). Combinatorial problems  Wikipedia (2025) involve finding a valid sequence of actions from an initial state $s_{0}$ to a goal state $s_{g}$, given a finite set of possible actions. Our framework, built upon the Graph-of-Thought inference approach  Besta et al. (2023), decomposes reasoning into discrete steps, each represented as a node in a structured inference graph. Through this inference graph, we simultaneously explore multiple chains of intermediate multimodal states toward the goal state. Below, we detail our method following the stages illustrated in Fig. 1. The full implementation, including prompts and outputs, is available in the paper’s GitHub repository.
Figure 3: Child state generation pipeline: the LMM selects an action from the parent node, generates a diagram schema and then an executable diagram code, performs self-reflection and verifies that action chosen does not violate any constraints and action path is feasible (example generations are from Blocksworld domain).
1. Generating Initial Conceptual Diagram for Domain.
See step 1 in Fig. 1. For each new planning domain, the LMM generates a domain-specific conceptual diagram that visually represents objects and their interactions using simple shapes, colors, and spatial arrangements. This is done entirely in a zero-shot manner, without manual engineering or external visual cues, conditioned only on the problem’s textual description. We begin by selecting a random instance from the domain and prompting the LMM to generate multiple candidate diagram schemas. The LMM then ranks these schemas based on how clearly they represent object positions and relationships, selecting the top-ranked one. Using this schema and the problem description, we generate multiple executable Matplotlib diagram codes to visualize the domain (Fig. 2). The LMM then ranks the rendered diagrams based on how effectively they visualize the structure and relationships between objects. The highest-ranked diagram is chosen as the ‘example’ domain-specific conceptual diagram, and its code serves as a reference for generating the diagram code for the initial and goal states of all instances in the domain.
2. Initial and Goal State Diagram Generation.
See step 2 in Fig. 1. We begin the inference process by generating diagrams for the initial state $s_{0}$ and the goal state $s_{g}$ of each instance, conditioned on the domain conceptual diagram code obtained in step 1. The LMM is instructed to strictly adhere to the representation of objects and their statuses as illustrated in the domain’s conceptual diagram, while accurately initializing the objects according to the specific instance in the diagram code.
3. Intermediate Child State Generation Pipeline.
See step 3 in Fig. 1. We denote by $s_{d}$ an intermediate state at depth $d$ of the graph, represented by a combination of: (i) textual description $T(s_{d})$; (ii) a diagram $D(s_{d})$; and (iii) the action path $A_{0:d}$ from the initial state $s_{0}$. We iteratively expand the inference graph depth by depth in a breadth-first search (BFS)  Cormen et al. (2009) manner, and apply beam search at each depth to select the top-k candidates for further expansion.
From each parent state $s_{d}$, we sample $n=4$ child states. W.l.o.g., next we describe the generation of a single such child state $s_{d+1}$ (Fig. 3). At each node, the LMM first selects the next candidate action $a_{d+1}$, conditioned on the parent state $s_{d}$ (which is represented by its diagram, textual description, and action path from initial state). We then generate the textual description $T(s_{d+1})$ of the resulting child state. The candidate action-state pair $(a_{d+1},T(s_{d+1}))$ is compared to previously generated child states to verify uniqueness. If different, the LMM generates a $s_{d+1}$ specific instance of the problem’s conceptual mapping $E(s_{d+1})$, which is a set of textual statements specifying shapes, sizes, positions, statuses (e.g., colors), and textual identifiers for each object in the state (see Fig. 3 for an example). The $E(s_{d+1})$ undergoes a self-reflection verification step to ensure consistency with $T(s_{d+1})$ and accuracy with respect to $a_{d+1}$, if failed we regenerate it. Subsequently, Matplotlib code $C(s_{d+1})$ is generated conditioned on $E(s_{d+1})$, $T(s_{d+1})$, and two example diagram codes: the initial state diagram code $C(s_{0})$ and the parent state diagram code $C(s_{d})$. Code is re-generated if it fails to run. After generating the child state diagram, we perform two additional action verification checks: (1) a local check that verifies the validity of the action $a_{d+1}$ based solely on the diagrams and textual descriptions of the parent and child states (ensuring the action complies with domain constraints); and (2) a global check that verifies whether the entire action path $A_{0:d+1}$ remains feasible for reaching the goal state $s_{g}$. If either of these checks fails, the child state is marked as invalid.
4. Inference Optimization: Beam Search.
See step 4 in Fig. 1. Applying Graph-of-Thought Besta et al. (2023) to combinatorial planning problems via naively expanding all nodes (e.g., BFS) would result in exponential growth of the search tree Russell and Norvig (2010). To mitigate this, we use a method inspired by beam search  Lowerre (1976), where first all states are expanded at each depth $d$ using BFS, generating up to $N$ child states per state. The LMM then ranks all candidate child states at depth $d+1$ based on their proximity to the goal state $s_{g}$, selecting only the top $k=4$ states for further expansion. This depth-wise ranking solves the exponential growth problem. Moreover, this ranking system resembles human problem-solving strategies, where shallow state-specific heuristics are employed to estimate how close intermediate states are to the goal  Gigerenzer and Goldstein (1996); Chronicle et al. (2006); Pizlo and Li (2005).
Figure 4: Sequence of intermediate state diagrams in the correct chain, from the initial state to the found goal state, for an example instance of five evaluated domains: Blocksworld, Floor Tiles, Elevator, Parking, and Tetris.
5. Depth-wise Backtracking.
See step 5 in Fig. 1. Additionally, we implement a depth-wise backtracking mechanism. If all candidate child states at a given depth fail verification, we backtrack to the deepest available ancestor nodes at depth $d_{\text{max}}$ and attempt new expansions. We allow up to $B=2$ backtracking attempts to any given depth. If all $B$ attempts at depth $d_{\text{max}}$ fail, we mark nodes at that depth as invalid and backtrack further to the next deepest available nodes at depth $d_{\text{max2}}$, where $d_{\text{max2}}<d_{\text{max}}$.
6. Search Termination.
See step 6 in Fig. 1. The inference process continues iteratively, expanding nodes depth-by-depth, until either the goal state $s_{g}$ is reached or a predefined computational budget is exhausted. Specifically, we set two types of computational limits: (1) a maximum number of generated states (120 states for simpler Blocksworld instances, 450 states for more complex domains), and (2) a maximum depth limit (28 for simpler Blocksworld instances, 100 for other domains). These limits ensure computational efficiency in terms of inference time and API usage costs. If the goal state is not reached within these constraints, the search is marked as incomplete. The goal state verification is performed by the LMM, which compares the diagram and textual description of the current state against the description of the goal state and its diagram (generated in (3) above).
In summary, our method leverages structured visual reasoning, self-generated conceptual diagrams, and optimized graph-based inference strategies to efficiently solve combinatorial planning problems with an LMM. Each step of our pipeline is visually illustrated in Figures 12, and 3.
4 Evaluation
To evaluate the proposed approach, we conducted experiments on seven different problem domains, including the popular Blocksworld (simple) Gupta and Nau (1991) and Blocksworld (hard) Valmeekam et al. (2024), as well as 5 new domains contributed in this work, prepared using a popular repository of PDDL problem instance generators ipc . Plan correctness was determined with VAL Howey et al. (2004). Experiments were run on a machine with dual-socket Intel Xeon Gold 5220R processors at 2.2 GHz, 35.75 MB L3, 48 cores per node, 8 nodes total).
Baselines.
We evaluate our approach against both single-inference and search-based inference methods. For single-inference baselines, we evaluate GPT-4o, o1-preview OpenAI (2024c), and o1-mini OpenAI (2024b), each prompted directly with PDDL instances using templates adapted from Valmeekam et al. (2023a). For search-based baselines, we evaluate Graph-of-Thought (GoT), which performs text-only breadth-first search over LLM-generated states, and Optimized GoT, which integrates beam search ($k=4$) into GoT to explore deeper solutions without exhausting the compute budget. We also compare against RAP Hao et al. (2023), which uses Monte Carlo Tree Search (MCTS) with the LLM generating both the world model and the plan. RAP’s compute budget is specified by iteration count; we adopted RAP(20) which was the highest compute budget authors used in their experiments. Although RAP was originally evaluated with LLaMA models, we extended it to use GPT-4o for direct comparison. However, RAP’s implementation requires hard-coded prompts and domain-specific parsers that are only available for Blocksworld, limiting its applicability to the other domains we evaluate. Together, these baselines probe the limits of purely text-based search and single-shot inference and provide strong points of comparison for our diagram-based framework.
Analysis Corrects (Accuracy) Incorrects Incompletes Avg Depth Max Depth Min Depth Avg Num States
Blocksworld (hard) 78% 10% 12% 20.15 36 18 177.7
Floor Tiles 36% 24% 40% 15.88 25 10 244
Parking 52% 40% 8% 9.6 29 2 149.1
Tetris 38% 60% 2% 6.45 11 4 46.24
Elevator 48% 42% 10% 20.89 26 16 160.30
Barman 30% 40% 30% 24.0 27 22 210.54
Table 2: Analysis of domains. ‘Corrects’: % instances with correct plans; ‘Incorrects’: % with incorrect plans; ‘Incompletes’: % terminated due to state budget; ‘Avg/Max/Min Depth’: avg/max/min number of actions in correct solutions; ‘Avg Num States’: average number of states generated across all instances of the domain.
Evaluated Domains.
We evaluate our method on combinatorial planning problems Ghallab et al. (2004) from the International Planning Competition (IPC) ipc (1998–Present), expressed in PDDL format McDermott et al. (1998). These problems are mostly NP-hard Gupta and Nau (1991). We evaluated our method on seven domains, namely Blocksworld (simple) and Blocksworld (hard) Gupta and Nau (1991); Valmeekam et al. (2024), and 5 other IPC domains: Floor Tiles Helmert (2000), Parking, Tetris Vallati (2015), Elevator, and Barman. For each domain, we generated instances using publicly available instance generators ipc . The Blocksworld domain involves stacking and unstacking blocks. We used the 500 simple instances in PlanBench Valmeekam et al. (2023a) with 3-5 blocks and 50 additional hard instances with 10-20 blocks (same instance complexity as in Valmeekam et al. (2024)). The Floor Tiles domain features robots painting tiles on a grid. We generated 50 PDDL instances for this domain with varying grid sizes (2-3 rows, 3-5 columns) and 1-2 robots. The Parking domain involves rearranging cars in curbs, with 50 instances using 4-5 curbs and 4-6 cars. The Tetris domain requires rearranging Tetris tiles on a grid, with 50 instances using $(4\times 4)$ or $(6\times 6)$ grids. Lastly, the Elevator domain simulates passenger transport in buildings, with 50 instances using 4-5 floors and 10-12 passengers. Refer to Fig. 4 for a sequence of intermediate state diagrams in the correct chain for an example instance of a subset of evaluated domains.
Translating PDDL to Natural Language and Back.
We first translate PDDL domains and instances into natural language. For domain translation, we use a manually engineered five-shot prompt containing translations of five different PDDL domains into natural language. For instance translation, we use a one-shot prompt containing a manually translated instance from the same domain. The pipeline operates on this text. After solving the problem using our pipeline, we translate the resulting natural language action sequence back into PDDL using a one-shot prompt of a random (incorrect) plan with accurate PDDL syntax.
4.1 Results and Analysis
Our main results are summarized in Table 1 which compares the performance of our method against leading models at the time of writing, including the specialized reasoning models o1-preview and o1-mini. Our approach significantly improves the performance of the base GPT-4o model and consistently outperforms strong search-based reasoning baselines across all evaluated domains. On the simpler Blocksworld instances, our method achieves 90.2% accuracy (+54.7 percentage points) and substantially outperforms all baselines, including text-only search-based methods like GoT (50%), Optimized GoT (58%), and RAP (58%), except for a slight disadvantage compared to the leading reasoning model, o1-preview (97.8%). We conjecture two reasons for o1-preview’s advantage: (i) the smaller problem size and shallow solution depth of simpler Blocksworld instances, and (ii) this planning domain is widely used, with many solved instances publicly available online, likely providing extensive exposure to o1 during its training.
On the much harder and deeper Blocksworld (hard) instances and the five new PDDL domains derived from the IPC competition ipc (1998–Present), where new instances are generated and solutions are not publicly available, our method demonstrates superior performance and scalability, supporting the conjectures above. For instance, our approach achieves 78% accuracy on Blocksworld (hard), while o1-preview’s performance drops to 23.65% and RAP’s to just 4%. This trend continues across all new domains, where our method consistently outperforms all baselines, including o1-preview. Standard GoT often fails entirely (e.g., 0% in Floor Tiles and Tetris) due to the combinatorial explosion of states, which exhausts the computational budget. While Optimized GoT improves on this with beam search, its performance still lags significantly behind our visual approach, highlighting that diagrams enable the model to more efficiently capture and analyze complex relational information that is difficult to parse from linear text alone.
Table 2 provides further insights into our method’s performance on these challenging domains. Despite significantly deeper solution paths (with average depths up to 24 actions per instance and up to 244 states generated per instance), our method successfully generates correct plans with as many as 36 sequential actions (Blocksworld (hard)). In contrast, GPT-4o and other search-based methods achieve very low performance (0-4%), highlighting its susceptibility to compounding errors in long-horizon compositional reasoning Valmeekam et al. (2024); Dziri et al. (2023). Our structured visual reasoning approach, incorporating backtracking and verification steps, effectively mitigates these compounding errors, enabling more robust performance even in deep and complex planning problems.
The primary limitation of our method observed in these experiments is the number of incomplete searches (e.g., 40% incomplete in Floor Tiles, 30% in Barman) which arise due to the computational budget constraint. Additionally, we note the method exhibits the highest incorrect rate (60%) on Tetris domain, primarily due to its inherently high branching factor (up to 24 possible actions per state) and complex action parameterization—each action can require up to 7 parameters detailing positions of all sub-tiles, compared to simpler domains like Blocksworld, where actions typically require only 1-2 parameters. Conversely, among the new domains evaluated, our largest performance gain over o1-preview occurs in the Barman domain (30% vs. 10%), likely due to our diagrammatic representation efficiently capturing high number of object interactions, which are challenging to represent for purely textual reasoning models.
Model Generalization.
To assess the generalizability of our framework, we evaluated it on other state-of-the-art LMMs using 50 instances from the Blocksworld (simple) domain. With LLaMA 4, our method increased accuracy from a baseline of 10% (single inference) to 74%, an over 7x improvement in accuracy. The improvement was also observed using Claude 3.5 Sonnet, where accuracy increased from 54.8% (using zero shot single inference as reported in Valmeekam et al. (2024)) to 98%, achieving state-of-the-art performance on this benchmark. These substantial gains demonstrate that the benefits of our approach are not tied to a specific model architecture but stem from the fundamental advantage of using model-generated conceptual diagrams as a reasoning medium.
Runtime and Cost Analysis.
We analyzed the runtime and API costs of our method and all other search-based baselines on 20 instances per domain. On the Blocksworld (simple) domain, our method had a median runtime of 381 ($\sim$6 minutes) seconds and a median cost of \$1.04 per instance. For more complex domains, the median runtime was 1038 seconds ($\sim$17 minutes) with median cost of \$2.98 per instance. Our approach is significantly more efficient than other search-based methods. On average, it was 31% faster and 36% cheaper than the GoT baseline, and 46% faster and 52% cheaper than RAP20+GPT-4o across all domains, while achieving substantially higher accuracy. Compared to the text-only Optimized GoT, incorporating diagrams added 213 seconds in latency and \$0.71 on average, measured across all domains, but this overhead yielded a 30 percentage point accuracy gain.
Ablation Corrects (Accuracy) Incorrects Incompletes Avg Depth Max Depth Avg Num States
GPT-4o + Visual Thinking 90.2% 6.4% 3.4% 10.07 28 38.89
No Diagram (Optimized GoT) 58% 36% 6% 8.28 18 25.86
No Diagram Schema 72% 20% 8% 7.41 18 46.2
No Code Execution 24% 66% 10% 7.82 22 32.7
1-Branching Factor 52% 4% 44% 6.38 16 23.9
2-Branching Factor 70% 6% 24% 8.29 24 26.27
No Backtracking 62% 4% 34% 6.06 14 22.12
No Beam Search 72% 4% 24% 6.61 12 58.31
Table 3: Ablation Study Results. ‘Corrects’ is our main accuracy metric. See Tab. 2 for columns notation.
4.2 Ablation Studies
To systematically evaluate the contributions of different components of our proposed framework, we conducted ablation studies on 50 instances from the Blocksworld domain. We examined the impact of state diagrams, diagram schema, diagram code execution, different branching factors, and inference optimizations (beam search and backtracking). Table 3 summarizes the results of these experiments.
Impact of Various Components of State Diagram Generation.
We first assessed the effectiveness of state diagrams by removing them entirely from the inference pipeline ("No Diagram"), resulting in a Graph-of-Thought approach that operates in text-only, rather than a multimodal approach. This led to a substantial performance drop from 90.2% to 58%, highlighting the critical role diagrams play in succinctly representing spatial and relational information. Additionally, the average solution depth of correctly solved instances decreased from 10.07 to 8.28, indicating that without diagrams, the model struggled to solve more complex instances requiring deeper reasoning. In another experiment, we removed the diagram schema ("No Diagram Schema") from the child state generation pipeline, directly inferring diagram code from the textual description of the state. This caused performance to drop from 90.2% to 72%. Although less severe than removing diagrams entirely, this drop demonstrates that diagram schemas effectively capture important spatial information from the textual description, aiding more accurate diagram generation.
Finally, we tested removing the rendered diagrams, providing only the Matplotlib code ("No Code Execution") when generating the next action. This resulted in the most significant performance drop (90.2% to 24%), even below the GPT-4o baseline (35.5%), clearly demonstrating that even though the code encodes the same spatial information as the diagram, the way this information and the interdependencies between objects are represented is crucial for model performance. Using the diagram code directly distracts the model and impairs reasoning, aligning with prior findings that extraneous details negatively impact model performance Shi et al. (2023); Liu et al. (2023b). These results reinforce the importance of diagrams as compact, intuitive representations that facilitate rapid perceptual inference and clear spatial reasoning Larkin and Simon (1987); Tversky (2011); Hegarty (2004), and that it is the representation of information in a multi-dimensional (2D, color-encoded) format that significantly aids spatial understanding and inference in models.
Impact of Branching Factor.
We next investigated the effect of branching factor of the inference graph (the number of candidate child states generated per state) on performance. Reducing the branching factor from 4 (our default) to 2 ("2-Branching Factor") decreased performance from 90.2% to 70%, primarily due to a significant increase in incomplete searches (24% vs. 3.4%). This suggests that exploring the third or fourth candidate child states, generated at higher temperatures, is frequently necessary to find the correct solution path. Thus, reducing branching factor limits randomness and diversity in candidate generation, leading to more incomplete searches. Additionally, the average depth of correctly solved instances decreased from 10.07 to 8.29, indicating difficulty solving problems with deeper solution depths with fewer candidate states. Further reducing the branching factor to 1 ("1-Branching Factor"), effectively converting our graph-based inference into a multimodal chain-of-thought approach (with diagrams), caused performance to drop even further to 52%. Despite this significant drop, performance remained substantially higher than the GPT-4 baseline (36%), underscoring the value of diagrams in improving reasoning performance of the LMM even without extensive search.
Impact of Inference Optimizations (Backtracking and Beam Search).
Finally, we evaluated the importance of our inference optimizations (backtracking and beam search) on top of the Graph-of-Thought (with diagrams) approach. Removing backtracking ("No Backtracking")—yielding a tree-of-thought method with beam search—reduced performance from 90.2% to 62%, primarily due to a rise in incomplete searches (34% vs. 3.4%). This increase arises because the LMM verification steps occasionally produce false negatives, incorrectly invalidating correct states. Without backtracking, the model cannot recover from these errors, leading to incomplete searches as no validated nodes remain at the frontier of the inference graph to expand from. Similarly, removing beam search ("No Beam Search") dropped performance to 72%, with rate of incomplete searches increasing to 24% . However, unlike backtracking, incompletes here result from exponential growth in the search space, causing the model to exhaust its computational budget (maximum allowed number of states to generate) before reaching the goal state. Indeed, the average number of states generated increased significantly (58.31 vs. 38.89), highlighting beam search’s critical role in managing combinatorial explosion. Both optimizations proved essential for solving deeper combinatorial problems, as shown by the reduced average solution depth without them (6.61 without beam search, 6.06 without backtracking, vs. 10.07 with both). These results demonstrate that backtracking and beam search are complementary and crucial optimizations for effective graph-based combinatorial planning.
5 Conclusion
Contributions.
In this paper, we introduced Visual Thinking, a zero-shot framework enabling LMMs to reason using text and self-generated conceptual diagrams to solve combinatorial planning problems. Our key contributions include: (1) autonomously generating conceptual diagrams directly from textual descriptions; (2) inference as a visual Graph-of-Thought with multiple chains of intermediate textual and diagrammatic representations; (3) employing beam search and depth-wise backtracking for efficient search; and (4) significantly outperforming GPT-4o and the specialized reasoning model o1-preview across multiple challenging planning domains.
Limitations.
As with any search-based inference method, our framework incurs additional computational cost and inference time by exploring multiple chains of thought and including visual representations for each state. Moreover, while setting a larger branching factor or beam width for the search tree can improve performance, it leads to higher computational cost and latency. However, we note that our method proved more computationally efficient than prior text-only search approaches like standard Graph-of-Thought (GoT) and RAP.
Future Work and Real-World Applications.
This work demonstrates a path for LMMs to move beyond purely textual reasoning toward a more powerful, human-like process that integrates reasoning through visual abstractions. Future work can extend this framework from combinatorial planning to more abstract domains, enabling LMMs to generate multimodal outputs such as software architecture diagrams, figures to visualize scientific hypotheses and causal dependencies, or tailored visual aids for educational contexts.
Our approach also has immediate applications in real-world combinatorial problems governed by flexible, natural-language constraints that challenge traditional algorithms. For instance, it could optimize warehouse layouts or product planograms based on qualitative rules (e.g., “group seasonal items together”) or resolve complex scheduling problems involving nuanced user preferences. In these contexts, the diagrams serve a dual role: providing a structured world model for the model to reason through and an intuitive validation tool for humans. This ability to reason and communicate multimodally is crucial for building AI that can tackle complex scientific, creative, and planning tasks.